OpenStudy (anonymous):

what is the derivative of y= 1/(2sin2x)

OpenStudy (anonymous):

try using the quotient rule or the chain rule

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2sin(2x)}\implies \cfrac{1}{2}\cdot [sin(2x)]^{-1}\) <-- chain rule it

OpenStudy (anonymous):

so it would be - csc2x cot2x?

OpenStudy (jdoe0001):

I got \(\bf \cfrac{1}{2sin(2x)}\implies \cfrac{1}{2}\cdot [sin(2x)]^{-1} \\ \quad \\ \quad \\ -\cfrac{1}{2}\cdot [cos(2x)]^{-2}\cdot 2\)

OpenStudy (anonymous):

i thought since the sin 2x would turn into csc 2x because thats the inverse of sin

OpenStudy (jdoe0001):

hmm well, the notation for the inverse is \(\bf ^{-1}\) yes however what we're taking is not the inverse function, but the reciprocal

OpenStudy (anonymous):

ok so with the answer you got, it would translate to \[-\csc^22x \]